OpenStudy (anonymous):

2^(x+2)=6^(2x-5) I do not know how to solve this problem. Help is greatly appreciated...thanks.

OpenStudy (anonymous):

solve for x... x+2=0 x=-2 2x-5=0 2x=5(divide both sides by 2 to eliminate 2) x=5/2 substitute values of x.. 2^(-2+2)=6^(2(5/2)-5) 1=1 :D

OpenStudy (anonymous):

You will need to use the log law: \[\log(a^{x}) = xlog(a)\] You can take the log of both sides of the equation and then you can solve for x